### v1

# Localization behavior in a Hermitian and non-Hermitian Raman lattice

Abstract
We introduce a flexible Raman lattice system for alakaline-earth like atoms to investigate localization behaviors in a quasi-periodic lattice with controllable non-Hermiticity. We demonstrate that critical phases and mobility edges can arise by adjusting spin-dependence of the incommensurate potentials in the Hermitian regime. With non-Hermiticity introduced by spin-selective atom loss, we find the localization behaviour in this system can be suppressed by dissipation. Our work provides insights into interplay between quasi-periodicity and non-Hermitian physics, offering a new perspective on localization phenomena.
I I. Introduction
In the presence of a disorder potential, quantum systems can exhibit exponential localization of wave functions induced by wave interference, a phenomenon known as Anderson localization Anderson (1958). Ultracold atom systems, with their exceptional purity and high controllability, enable detailed exploration of localization phenomena. Recent experiments have investigated Anderson localization in optical lattices with either random disorder Billy et al. (2008) or incommensurate quasi-periodic structures Roati et al. (2008); Lüschen et al. (2018); Xiao et al. (2021).
For ultracold atoms, quasi-periodic optical lattices are created by superimposing two optical lattices with incommensurate wavelengths. In such setups, a notable transition from extended to localized states, commonly referred to as the Anderson transition, can manifest even within one-dimensional configurations Grempel et al. (1982); Aulbach et al. (2004), contrary to uncorrelated disordered systems where such transitions are expected to occur only in dimensions higher than two according to the scaling theory Thouless (1974); Abrahams et al. (1979). These systems also allow for the engineering of mobility edges-boundaries that separate coexisting localized and extended states at discrete energy levels Sarma et al. (1988); Evers and Mirlin (2008); Biddle and Das Sarma (2010); Lüschen et al. (2018); An et al. (2018, 2021). Recent theoretical studies suggest even more complex behavior, including intermediate critical phases and hybrid phases where localized, extended, and critical states coexist in specific one-dimensional quasi-periodic models Wang et al. (2020, 2022).
Recent years have seen growing interest in non-Hermitian physics El-Ganainy et al. (2018); Ashida et al. (2021) across multiple research domains, including electrical circuits Helbig et al. (2020); Ezawa (2019), optics and photonics Lee et al. (2009); Xiao et al. (2020); Zhou et al. (2018); Weidemann et al. (2020), optomechanics Patil et al. (2022), NV centers Zhang et al. (2021); Yu et al. (2022), trapped ions Cao et al. (2023); Zhang et al. (2025), superconducting qubits Chen et al. (2021) and ultracold atom systems Li et al. (2019); Liang et al. (2022); Ren et al. (2022). Non-Hermitian systems display unique features such as exceptional points or rings Zhen et al. (2015); Ren et al. (2022); Liu et al. (2025), novel topological invariants, and the non-Hermitian skin effect Liang et al. (2022); Zhao et al. (2025). The Hatano-Nelson model, a fundamental non-Hermitian topological model, combines asymmetric hopping amplitudes with an on-site random potential Hatano and Nelson (1996). However, a comprehensive understanding of the interplay between disorder and non-Hermiticity, particularly the influences of dissipation on localization behavior, remains largely unexplored Zeng et al. (2017); Zhang et al. (2020); Harter et al. (2016). Furthermore, experimental methods for studying this relationship between disorder and non-Hermiticity remain limited.
In this paper, we introduce a flexible Raman lattice scheme Liu et al. (2013) tailored for 173Yb atoms Song et al. (2018, 2019) or alkaline-earth metal atoms Liang et al. (2023) to investigate localization behaviors. The Raman lattice configuration induces quasi-periodicity using two optical lattices with incommensurate wavelengths, with the weak lattice potential akin to an incommensurate Zeeman potential. By adjusting the frequency of the incommensurate lattice close to the ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition, the spin-dependent nature of the lattice potential may lead to the emergence of critical phases or mobility edges within the system. Finally, we generalize the Raman lattice with incommensurate potentials into the non-Hermitian regime with dissipation induced by atom loss Zhao et al. (2025). The analysis reveals that dissipation exerts a significant influence on the localization behaviors within the system, offering a platform to investigate the intricate interplay between disorder potential and non-Hermitian physics.
The subsequent sections of this paper are structured as follows: Sec.II presents the model Hamiltonian of the flexible Raman lattice and outlines its experimental realization scheme. In Sec.III-V, we delve into the impact of the spin-dependency of the incommensurate lattice and dissipation within the system. Our focus will encompass three distinct scenarios: a completely spin-dependent incommensurate lattice, a partially spin-dependent incommensurate lattice, and an incommensurate lattice coupled with dissipation. Finally, Sec.VI offers a concise summary of the findings and conclusions derived from this investigation.
II II. Model Hamiltonian
The Hamiltonian considered in this paper can be written as:
$H=H_{0}+\sum_{i,\sigma=\uparrow,\downarrow}(\delta_{\sigma}^{i}+i\gamma_{\sigma})n_{i,\sigma}$ (1)
where $\delta^{i}_{\sigma}=M_{z,\sigma}\cos{\left(2\pi\beta i\right)}$, with $M_{z,\sigma}\propto V_{s,\sigma}$ and $\gamma_{\sigma}$ denotes the incommensurate Zeeman potential strength and dissipation strength for spin $\sigma=\uparrow,\downarrow$, respectively. $n_{i,\sigma}=c^{\dagger}_{i,\sigma}c_{i,\sigma}$ is the particle number operator, $\beta$ is an irrational number and:
$H_{0}$ $=-t_{0}\sum_{i,\pm,\sigma}c_{i,\sigma}^{\dagger}c_{i\pm 1,\sigma}$ (2)
[…]
Here, $t_{0}$ ($t_{so}$) represents the spin-conserved (spin-flip) hopping strength between neighboring sites. The Hamiltonian renders a 1D AIII class topological insulator if replacing the incommensurate Zeeman potential with a constant uniform Zeeman potential Song et al. (2018); Liu et al. (2013); Wang et al. (2018); Zhang and Liu (2018).
Figure 1: Experimental realization of critical phase in optical Raman lattice with 173Yb atoms. (a) Experimental setup of the optical Raman lattice consists of a standard 1D optical Raman lattice $E_{1}$, $E_{3}$ with another spin-dependent incommensurate lattice $E_{21}$ and $E_{22}$. (b) Schematic energy diagram with relevant transitions. (c) The incommensurate lattice induces a spin-dependent offset $V_{2,\sigma}(x)$ to the standard optical Raman lattice potential. (d) For the implementation of a perfect spin-dependent incommensurate lattice, the laser detuning should be set approximately 1.9 GHz blue-detuned from the ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition.
This Hamiltonian can be experimentally implemented through the configuration outlined in Fig. 1(a). To establish the conventional optical Raman lattice potential Liu et al. (2013); Wu et al. (2016); Song et al. (2018), a standing-wave beam $E_{1}$ with $y$ polarization and a free-running beam $E_{3}$ with z polarization are employed. The single-photon detuning of these two beams can be finely tuned to approximately 11 GHz blue-detuned from the ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition, as depicted in Fig. 1(b). In this setup, the primary lattice, characterized by a wave vector of $k_{0}=2\pi/\lambda$, exhibits nearly spin-independent lattice potential $V_{1,\sigma}(x)=V_{p,\sigma}\cos^{2}{k_{0}x}$ with $V_{p,\uparrow}/V_{p,\downarrow}\sim 1.01$, where $V_{p,\uparrow}/V_{p,\downarrow}$ denotes the primary lattice depth for spin up or spin down atoms. In the subsequent discussion, we will neglect the spin-dependent effects of the primary lattice potential ($V_{p,\uparrow}=V_{p,\downarrow}=V_{p}$). Moreover, two additional beams propagating at an angle of 52 degrees relative to the x direction are utilized to induce a spin-dependent incommensurate lattice featuring a wave vector of $k_{1}=k_{0}\cos{\theta}$, giving rise to a secondary incommensurate Zeeman potential $V_{2,\sigma}=V_{s,\sigma}\cos^{2}{k_{1}x}$ with $V_{s,\sigma}$ representing the incommensurate lattice depth, as illustrated in Fig. 1(c). The spin-dependence of this incommensurate lattice can be controlled through the detuning $\Delta$ of this lattice beam with respect to the ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition. When $\Delta\sim 1.9~{}\text{GHz}$, the incommensurate lattice becomes entirely spin-dependent with $M_{z,\uparrow}/M_{z,\downarrow}=-1$ (Fig. 1(d)). To achieve a complete spin-independent configuration, the detuning of the incommensurate lattice can be set larger than 13GHz. The Raman coupling in this setup can be denoted by $\mathcal{M}(x)=M_{R}\cos{k_{0}x}$. The introduction of dissipation into this Raman lattice setup can be realized through adding a nearly resonant loss beam with wavelength close to ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition Zhao et al. (2025).
For simplicity, all phases along the y direction in these potentials are neglected during the calculation and the Hamiltonian governing this experimental setup can be expressed as:
$H$ $=\left[\frac{\hbar^{2}k_{x}^{2}}{2m}+V_{1}(x)\right]\otimes\mathds{1}+\sum_{\sigma=\uparrow,\downarrow}V_{2,\sigma}(x)|\sigma\rangle\langle\sigma|$ (3)
$+\mathcal{M}(x)\sigma_{x}+\frac{\delta}{2}\sigma_{z}+\sum_{\sigma=\uparrow,\downarrow}\frac{i}{2}\gamma_{\sigma}|\sigma\rangle\langle\sigma|$
[…]
III III. Localization behavior in the Hermitian regime
Except for the localized phase, recent theoretical investigations have predicted the existence of a third fundamental phase, called critical phases, positioned between the localized and extended phases within such a spin-orbit coupled optical lattice featuring an incommensurate Zeeman potential Wang et al. (2020); Song et al. (2018); Zhao et al. (2025). Noteworthy features of these critical phases encompass critical spectral statistics Geisel et al. (1991); Machida and Fujita (1986); Bertrand and García-García (2016), multifractal properties of wave functions Halsey et al. (1986); Mirlin et al. (2006); Dubertrand et al. (2014), and dynamical evolutions Hisashi and Shuji (1988); Ketzmerick et al. (1997); Larcher et al. (2009). Moreover, recent theoretical investigations have also unveiled the possibility of a quantum phase harboring coexisting localized, extended, and critical regions within a one-dimensional quasi-periodic model Wang et al. (2022).
Therefore, we first consider an entirely spin-dependent incommensurate lattice ($M_{z,\uparrow}/M_{z,\downarrow}=-1$) in the Hermitian regime ($\gamma_{\sigma}=0$), and find that a critical phase manifests in this configuration. The phase of this system can be characterized through the mean fractal dimension $\bar{\eta}$, as introduced by Wang et al. (2020):
$\bar{\eta}=-\frac{\lim_{L\rightarrow\infty}\ln{\left[(2L)^{-1}\sum^{2L}_{m=1}\sum^{L}_{j=1}\left(u^{4}_{m,j}+v^{4}_{m,j}\right)\right]}}{\ln{2L}}$ (4)
Here, $\bar{\eta}$ is averaged over all eigenstates, which can be represented as:
$|{\psi_{m}}\rangle=\sum_{j=1}^{L}\left(u_{m,j}c^{\dagger}_{j,\uparrow}+v_{m,j}c^{\dagger}_{j,\downarrow}\right)|{\text{vac}}\rangle$ (5)
It is established that, in extended (localized) states, the value of $\bar{\eta}$ tends towards 1 (0), while within the critical phase, $10<\bar{\eta}<$ 1.
In Fig.2(a), the mean fractal dimension is depicted as a function of $M_{z}$ and $t_{so}$, revealing a distinct region where $10<\bar{\eta}<$ 1, indicating the emergence of the critical phase. The phase boundary can be delineated by $M_{z}=2|t_{0}\pm t_{so}|$, where the symbol $+$ signifies the boundary between localization and critical phases, and -- signifies the boundary between extended and critical phases Wang et al. (2020). Fig. 2(b) shows the fractal dimension for various eigenenergies and $M_{z}$ when $t_{so}=0.3$, illustrating that all individual states within the critical region exhibit critical behavior. However, if the incommensurate lattice is entirely spin-independent ($M_{z,\uparrow}/M_{z,\downarrow}=1)$, the critical phase region disappears, as illustrated in Fig. 2(c). Although a region where $\bar{\eta}$ deviates significantly from 1 and 0 can be identified, a further examination of the fractal dimension $\eta$ of the corresponding eigenstates as a function of $M_{z}$ suggests an intermediate regime characterized by the coexistence of extended and localized states rather than a critical phase (Fig.2(d)).
Figure 2: Phase diagram of the optical Raman lattice system with additional incommensurate lattice. (a) The various phases, namely extended, critical, and localization phases, are distinguished based on the mean fractal dimension $\bar{\eta}$ when $M_{z,\uparrow}/M_{z,\downarrow}=-1$. $M_{z}$ and $t_{so}$ are presented in units of $t_{0}$. (b) Fractal dimension $\eta$ and energy of individual states for distinct values of $M_{z}$ with $t_{so}=0.3$ and $M_{z,\uparrow}/M_{z,\downarrow}=-1$. (c) Phase diagram of mean fractal dimension $\bar{\eta}$ as a function of $M_{z}$ and $t_{so}$ when $M_{z,\uparrow}/M_{z,\downarrow}=1$. Other parameters are identical to those in b. (d) Fractal dimension $\eta$ and energy of individual states for distinct values of $M_{z}$ with $t_{so}=0.3$ and $M_{z,\uparrow}/M_{z,\downarrow}=1$. All results are simulated with $L=1597$ and $\beta=987/1597$ in periodic boundary condition.
In the pursuit of experimentally detecting localization behavior in an optical lattice system, one common method involves observing the expansion dynamics, as conducted in various studies Billy et al. (2008); Roati et al. (2008); Lüschen et al. (2018). Fig. 3(a) shows simulated time evolutions of wave packets positioned in the extended, critical, and localized phase regions. These simulations are based on an initial wave packet configuration defined as $|{\psi_{j}(t=0)}\rangle=(\sqrt{\pi}a)^{-1/2}e^{-(j-j_{0})^{2}/2a^{2}}|{\uparrow}\rangle$, where $a$ represents the half-width, and the packet is centered at site $j_{0}$. Notably, the outcomes reveal distinctive behaviors: in the extended phase, the atom cloud exhibits ballistic expansion, while in the localized phase, occupancy remains predominantly restricted to the initial site. In the critical phase, a quasi-localization phenomenon is observed, with the atom cloud spreading to neighboring unit cells around the initial site, without fully occupying the entire system.
To further characterize the expansion dynamics, the mean square displacement can be employed as an observable, defined as Wang et al. (2020):
$W(t)=\left[\sum_{j,\sigma}(j-j_{0})^{2}\langle n_{j,\sigma}(t)\rangle\right]^{1/2}$ (6)
In Fig.3(b), the evolution of $W$ over time for various $M_{z}$ values, with $a=5$, is presented. In the extended phase, $W$ rapidly converges to a large and stable value, while in the localized phase, $W$ remains consistently small. Conversely, in the critical phase, $W$ exhibits gradual and slow growth, distinguishing it from both the extended and localized regimes. Furthermore, the determination of phases can be facilitated by the time-averaged observable $\bar{W}=\frac{1}{N_{t}}\sum_{m=1}^{N_{t}}W(m\Delta T)$, where $N_{t}=100$ and $\Delta t=20$, representing the average width of the wave packet throughout the temporal evolution, as depicted in Fig. 3(c).
Figure 3: Simulation of detecting critical phase with expansion dynamics. (a) Time evolution of the wave packet for the system located in the extended phase ($M_{z}=0.5$), critical phase ($M_{z}=2.0$), and localized phase ($M_{z}=3.5$). (b) Expansion dynamics of the wave packet, characterized by $W$, as a function of $M_{z}$, with $t_{so}=0.3$ and an initial width of $a=5$. (c) Time-averaged $\bar{W}$ as a function of $M_{z}$ with different initial width $a$. All results are simulated with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary condition.
In most localization behavior studies, a crucial consideration lies in the necessity for the initial state to populate only a few lattice sites to ensure clarity in distinguishing between different phases (Fig.3(c)). While experimental setups in ultracold atom investigations enable the detection and preparation of Gaussian wavepackets as initial states through techniques like a matter wave magnifierMurthy et al. (2014); Asteria et al. (2021) or a quantum gas microscope Bakr et al. (2009); Sherson et al. (2010); Miranda et al. (2015); Haller et al. (2015); Edge et al. (2015); Omran et al. (2015); Cheuk et al. (2016); Parsons et al. (2016); Yamamoto et al. (2016); Mitra et al. (2018); Yamamoto et al. (2020); Kwon et al. (2022), this can introduce complexity and inconvenience to the experimental arrangement. An intriguing observation emerges from numerical simulations represented in Fig. 3(a), indicating that in the extended phase, the presence of spin-down atoms is negligible when all initial states are occupied by spin-up atoms. Contrastingly, in the critical phase, clear evolution of spin-down atoms is observable over time. This can provide another way to detect the emergence of a critical phase in the current system.
In Fig.4(a), the spin polarization for systems where all lattice sites are initially uniformly occupied by spin up atoms is depicted for various phases. Within the extended phase, spin polarization exhibits minor oscillations with the value close to 1, while in critical and localized states, distinct oscillations occur, with the critical phase displaying larger oscillation amplitudes than the localized phase. Comparing with the distinct boundary between the extended and critical phases, the demarcation between the critical and localized phases, while not as pronounced, remains identifiable. Utilizing time-averaged spin polarization can aid in characterizing the emergence of the critical phase (Fig.3(c)-(d)). Notably, the time-averaged spin polarization tends towards zero in the critical phase, contrasting with higher values outside this phase.
Figure 4: Evolution of spin polarization for different phases. (a) Time evolution of the spin polarization for the system located in the extended phase ($M_{z}=0.5$), critical phase ($M_{z}=2.0$), and localized phase ($M_{z}=3.5$) with the initial state comprising spin-up atoms in a uniform distribution. (b) Spin dynamics as a function of $M_{z}$, with $t_{so}=0.3$, and an initial distribution in the spin up state uniform distribution. (c) Time-averaged spin polarization as a function of $M_{z}$ averaged over $t=0-200\text{s}$, which show minimal disparity between initial states of uniform distribution or Gaussian distribution. (d) Time-averaged spin polarization as a function of $M_{z}$ and $t_{so}$. All simulations are conducted with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary conditions.
IV IV. Coexistence of different states
Returning to the spin-independent incommensurate lattice case, although the critical phase is absent in such a system, under specific conditions, extended and localized states can coexist, separated by a mobility edge. For a more general situation, where the incommensurate Zeeman potential exhibits a blend of spin-dependence and independence, a phase emerges encompassing coexisting localized, extended, and critical regions Wang et al. (2022). Illustrated in Fig.5(a) is the fractal dimension corresponding to various eigenenergies and $M_{z}$, revealing regions where three or two distinct types of eigenstates coexist, separated by mobility edges. To validate the coexistence of difference phases within this domain, we consider the case of $M_{z}=1.5$ as an example (refer to Fig.5(b)). Notably, as the system size increases, $\eta$ converges towards 0 for all states in zones I and IV, indicative of localization states, while tending towards 1 for all states in zone III, signifying extended states. Conversely, within zones II and V, the fractal dimension markedly deviates from 0 and 1 and remains relatively constant with system size, suggesting a critical state in these regions.
The time evolution of the wave packet under such circumstances, as depicted in Fig.5(c), showcases characteristics signatures of these three distinct phases concurrently. On the whole, the time evolution of the wave packet exhibits quasi-localization, highlighting the presence of critical states. However, the higher concentration of atom clouds in the vicinity of central sites, particularly for spin-down atoms, indicates the existence of localized states. Furthermore, the inset of Fig.5(c) illustrates instances of ballistic expansion behavior of localized states upon closer inspection of the color bar. Through simulations of time evolution, the coexistence of extended, critical, and localized states under such conditions is further substantiated.
Figure 5: Phase with coexisting localized, extended, and critical zones. (a) Fractal dimension $\eta$ and energy of individual states for distinct values of $M_{z}$ with $t_{so}=0.3$ and $M_{z,\uparrow}/M_{z,\downarrow}=-2$ with $t_{so}=0.3$. Simulations are conducted with $L=1597$ and $\beta=987/1597$ under periodic boundary condition. (b) Fractal dimension of different eigenstates as a function of eigenenregy for different system size $L=233$ and $L=1597$ with $M_{z}=1.5$ and $t_{so}=0.3$. For convenience, only the region with $E<0$ is presented , which is symmetric to that with $E>0$. (c) Time evolution of the wave packet for the system located in the phase with coexisting localized, extended, and critical zones ($M_{z}=1.5$, $t_{so}=0.3$ and $M_{z,\uparrow}/M_{z,\downarrow}=-2$). Simulations are conducted with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary condition.
V V. Suppression of critical phase in non-Hermitian Raman lattice
Recently, various non-Hermitian extensions of systems exhibiting Anderson localization have been explored, such as the non-reciprocal Aubry-André model Jiang et al. (2019) and the PT-symmetric Aubry-André model Yuce (2014); Liang et al. (2014); Hang et al. (2015); Harter et al. (2016); Zeng et al. (2017); Longhi (2019). Numerical studies have delved into the impact of nonreciprocal tunneling, physical gain and loss on parity-time symmetry breaking, the Hofstadter butterfly spectrum Longhi (2014); Yuce (2014); Liang et al. (2014); Harter et al. (2016), topological phase transitions Longhi (2019), and the localization characteristics of eigenstates Hatano and Nelson (1996); Shnerb and Nelson (1998); Hang et al. (2015); Zeng et al. (2017); Jiang et al. (2019); Zhou (2023). Given these insights, it becomes pertinent to investigate the effects of non-Hermiticity on our system.
Expanding our setup into the non-Hermitian regime involves introducing dissipation to one of the spin states ($\gamma_{\sigma}\neq 0$) through a nearly resonant loss beam Zhou (2023). We focus on the scenario with $M_{z,\uparrow}/M_{z,\downarrow}=-1$, which exhibits a critical phase in the Hermitian context. The phase diagram for the system with dissipation ($\gamma_{\downarrow}=-2$) is illustrated in Fig. 6(a), revealing a transition of critical phases in the Hermitian regime to extended or localized phases in non-Hermitian regime. Fig. 6(b) presents the mean fractal dimension $\bar{\eta}$ for varying dissipation values $\gamma_{\downarrow}$ with $t_{so}={1.0}$. The plateau characterizing the critical phase in the Hermitian setup vanishes upon the introduction of dissipation. For system with larger $M_{z}$, the mean fractal dimension decreases, indicative of a transformation of eigenstates into localized states, whereas for lower $M_{z}$ values, the mean fractal dimension rises, signifying a transition towards extended states.
Furthermore, Fig. 6(c) shows the fraction of eigenstates exhibiting different fractal dimensions, depicting a critical-extended transition. Employing a methodology akin to that in Fig. 5, it is observed that in the Hermitian regime, the fractal dimension remains stable, indicative of critical states in these regions. However, upon introducing dissipation ($\gamma_{\downarrow}=2$), $\eta$ converges towards 1 for all states when the system size increases, confirming the transformation of nearly all states into extended states and the suppression of the critical phase within the system due to the induced dissipation (Fig. 6(d)). This suppression of the critical phase is attributed to non-Hermiticity, as evidenced by the persistence of this effect even under conditions where $\gamma_{\uparrow}/\gamma_{\downarrow}=-1$.
Figure 6: Suppression of critical phase in non-Hermitian Raman lattice. (a) The phases diagram after inducing dissipation ($\gamma_{\downarrow}=-2$) to the system, distinguished based on the mean fractal dimension $\bar{\eta}$ when $M_{z,\uparrow}/M_{z,\downarrow}=-1$. $M_{z}$ and $t_{so}$ are presented in units of $t_{0}$. The dashed lines show the boundary between different phases in Hermitian regime. (b) Mean fractal dimension $\bar{\eta}$ for different dissipation $\gamma_{\downarrow}$ as a function of $M_{z}$ with $t_{so}={1.0}$. (c) Fraction of eigenstates with $\eta>0.8$, $\eta<0.3$ and $0.80.3<\eta<$ 0.8 respectively as a function of $\gamma_{\downarrow}$ with $t_{so}=1.0$ and $M_{z}=1.0$. (d) Fractal dimension of different eigenstates as a function of eigenenergy for different system size $L=233$ and $L=1597$ with $M_{z}=1.5$ and $t_{so}=0.3$. The top one is without dissipation and the bottom one is with dissipation. For convenience, only the region with $E<0$ is presented , which is symmetric to that with $E>0$. Unless otherwise specified, all simulations are conducted with $L=1597$ and $\beta=987/1597$ under periodic boundary condition.
VI VI. Summary
This study introduces a versatile Raman lattice configuration aimed at investigating the localization behavior for alkaline-earth-like atoms, exemplified by 173Yb. The experimental setup explores three distinct quasi-periodic Raman lattice configurations: an entirely spin-dependent incommensurate lattice in a Hermitian regime, a partially spin-dependent incommensurate lattice under Hermitian conditions, and a fully spin-dependent incommensurate lattice in a non-Hermitian regime. In the case of the entirely spin-dependent incommensurate lattice in the Hermitian regime, the emergence of a critical phase is observed, which can be detected by expansion dynamics and spin polarization evolution. Conversely, when the incommensurate lattice transitions into a combination of spin-dependent and spin-independent characteristics, a new phase with coexistence of extended, critical, and localized phases separated by mobility edges shows up. On the other hand, we also find that introducing dissipation into the system leads to the suppression of the critical phase due to non-Hermiticity within specific parameter spaces. This work aims to stimulate further experimental investigations into the interplay between the localization behavior and non-Hermiticity Zhou (2023); Jiang et al. (2019) in ultracold atoms.
VII ACKNOWLEDGMENTS
GBJ acknowledges support from the RGC through 16306922, 16302123, 16305024, N-HKUST636-22, and RFS2122-6S04.